OpenStudy (anonymous):

Calculator question: Let f(x)=ln(e^2x)/(x-1) for x>1. If g is the inverse of f, then g'(3)= (A) 2 (B) 1 (C) 0 (D)-1 (E)-2

OpenStudy (anonymous):

1

OpenStudy (anonymous):

the answer is -2, I just don't know how to get there

OpenStudy (jziggy):

f(x) =ln(e^2x)/(x-1) First I can put the 2x up front. f(x) = 2xln(e)/(x-1) Ln(e) = 1 f(x) = 2x/(x-1) y = 2x/(x-1) Now to get the inverse. x = 2y/(y-1) xy - x = 2y xy - 2y = x y(x-2) = x y = x/(x-2) f'(x) = [(x-2) - x]/(x-2)^2 f'(x) = -2/(x-2)^2 f'(3) = -2/(3-2)^2 f'(3) = -2/1^2 f'(3) = -2